# The adjoint state method for parametric definable optimization without smoothness or uniqueness

### 27 Mar 2026 v1

@@ Abstract
Definable parametric optimization problems with possibly nonsmooth objectives, inequality constraints, and non-unique primal and dual solutions admit an adjoint state formula under a mere qualification condition. The adjoint construction yields a selection of a conservative field for the value function, providing a computable first-order object without requiring differentiation of the solution mapping. Through examples, we show that even in smooth problems, the formal adjoint construction fails without conservativity or definability, illustrating the relevance of these concepts to grasp theoretical aspects of the method. This work provides a tool which can be directly combined with existing primal-dual solvers for a wide range of parametric optimization problems.
Key words and phrases. Adjoint state method, conservative fields, nonsmooth optimization, o-minimal geometry, parametric optimization, variational analysis.
2020 Mathematics Subject Classification. Primary 90C31; Secondary 49J53, 03C64, 65K05, 90C26.
@@ 1 Introduction
Parametric optimization problems are central in optimization [9], control [31], and engineering [10]. They also arise naturally in modern machine learning, where one seeks to tune hyperparameters, architectures, regularization levels, or embedded optimization layers through the optimization of an induced value function; see [35, 4]. In these situations, the quantity of interest often has the parametric form
@@
where the parameter $\theta$ ranges in $\mathbb{R}^{q}$, $F\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}$ is locally Lipschitz, the constraint functions $G\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{m}$ and $H\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{p}$ are continuously differentiable, and the inequality is understood coordinate-wise. A central difficulty in optimizing the value function $f$ lies in obtaining useful first-order information at a reasonable computational cost. In principle, if one denotes by
@@
a selection of minimizers, the value writes $f(\theta)=F(x^{*}(\theta),\theta)$, thus, if the mapping $\theta\mapsto x^{*}(\theta)$ were differentiable, the chain rule would give
@@
However, this approach raises both computational and conceptual difficulties. On the computational side, differentiating through a solver is generally expensive, while conceptually it relies on strong regularity assumptions, typically requiring an application of the implicit function theorem, which is often restrictive in practice.
The adjoint state method, introduced by Céa [11], provides a way to bypass the computation of the derivative of the solution mapping. It also avoids relying on the duality framework typically associated with convexity assumptions [37, 21]. Denoting $(x^{*}(\theta),\lambda^{*}(\theta),\mu^{*}(\theta))$ a solution of the KKT system of (1.1), the adjoint formula writes (see [11] for a formal derivation under equality constraints)
@@
Remarkably, at a negligible additional cost—see Remark 2.8()—associated with computing Lagrange multipliers, this formula does not involve the derivative of the primal-dual solution mapping. Yet, the theoretical validity of the analysis of Céa requires solutions of the KKT system to be unique and differentiable.
@@
The objective of this work is to analyze the adjoint formula, under simple and versatile assumptions, that applies to a broad class of nonsmooth optimization problems with inequality constraints and possibly nonsmooth or non-unique primal and dual solutions. To this end, we show that the expression (1.2) carries strong variational information whenever the Mangasarian–Fromovitz constraint qualification holds and all the mappings $F,G,H$ belong to a common o-minimal structure (for example, are semialgebraic). We also discuss the computational cost of our approach, comparing it with less sophisticated methods and connecting it to automatic differentiation.
@@
Our approach relies on two structural ingredients: definability and conservative calculus. Definability provides a natural framework that encompasses most optimization and machine learning models [8, 15], while offering powerful stability and regularity properties that are essential for our analysis. Conservative calculus, on the other hand, supplies generalized differential objects suited to nonsmooth calculus and has already led to strong applications in optimization and learning; see [8, 15, 34]. The use of conservative fields for the value function is essential here: the formal application of the adjoint formula in (1.2) produces artifacts, even for smooth data $F,G,H$ resulting in a smooth value function $f$, see the very simple Example 2.7(). This is very common in nonsmooth differential calculus. Nonetheless, the outputs of the adjoint formula (1.2) form a conservative field, in particular, artifacts only have a minor impact. We also point out that the natural connections between conservativity, differentiable programming, and algorithmic differentiation make our approach amenable to modern computational frameworks [7, 8].
@@
Although primarily theoretical, this work provides a rigorous foundation for the practical use of the adjoint state method in settings where smoothness or uniqueness of solutions cannot be guaranteed. Such situations arise in a wide range of applications, including robust optimization [5], uncertainty quantification [1], optimal design of structures and systems [1], parametric decomposition [24, 20], and discretized neural ODEs [12]. A detailed investigation of these applications is beyond the scope of the present paper.
To conclude, we emphasize that the definable assumption is quite sharp. Indeed, in the absence of definability, the adjoint formula may fail to carry variational meaning for the value function, even for locally Lipschitz, path differentiable and (point-wise) differentiable objective function with smooth constraints. We provide an example showing that some structural assumption is required, justifying the o-minimal (definable) framework to ensure the validity of the adjoint formula. This framework is broad enough to cover most models that arise in applications. For example, most relevant neural networks are definable in some o-minimal structure [8, 15]. Beyond machine learning, many engineering and numerical models, when discretized, involve solely polynomials or definable functions.
Related work.
Let us briefly describe some related literature. In the continuous, infinite dimensional setting, extensions of the adjoint method to nonsmooth or non-unique solutions can be found in [16, 43]. They typically provide results on directional derivatives and rely on saddle-point-type arguments of a $t$-parametrized Lagrangian. The saddle point argument was removed later in [17, 18, 19]. A naive adaptation of these previous works to our discrete setting would yield, at best, only directional derivatives, rather than a directly computable gradient-like object for the value function, as provided by our approach. Moreover, it would require, when possible, solving costly optimization problems over the sets of primal and dual solutions. In nonlinear programming, related sensitivity results based on Lagrange multipliers likewise provide only upper and lower bounds on directional derivatives and entail similar optimization procedures over both primal and dual solution sets; see the work of Gauvin [23] and subsequent extensions [28, 22, 33, 3, 13, 38, 41]. Beyond directional derivatives, [27] and [36, Theorem 10.13] established that the Clarke subdifferential of the value function is included in a convex set constructed with the subdifferential of the Lagrangian; see the discussion in Section 4.2. However, as noted in Example 2.7(), this inclusion is strict in general, further confirming that subdifferentials are not the best paradigms to analyze the adjoint state method with multiple solutions. A limitation that this paper overcomes through conservative gradients.
@@
Let $\ell,L\in\mathbb{N}$ be generic natural numbers. We denote by $\langle{\cdot,\cdot}\rangle$ the canonical Euclidean scalar product on $\mathbb{R}^{\ell}$ and $\lVert{\cdot}\rVert$ its associated norm. For $r>0$ and $x\in\mathbb{R}^{\ell}$, we denote by $\mathcal{B}_{c}(x,r)$ the closed ball centered at $x$ with radius $r$, and by $\mathcal{B}(x,r)$ its open counterpart. Let $A\subset\mathbb{R}^{\ell}$. We write $\operatorname*{cl}A$, $\operatorname*{int}A$, and $\operatorname*{conv\!}A$ for the closure, interior, and convex hull of $A$, respectively. The Jacobian and partial Jacobian of a vector-valued function $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}^{L}$ at $z=(x,y)\in\mathbb{R}^{\ell}$ are denoted by $\operatorname{Jac}\varphi(z)$ and $\operatorname{Jac}_{x}\!\varphi(x,y)$, respectively. In the same way, the gradient and partial gradient of a real-valued function $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}$ at $z=(x,y)\in\mathbb{R}^{\ell}$ are written as $\nabla\varphi(z)$ and $\nabla_{\!x}\varphi(x,y)$, respectively. For $s,t\in\mathbb{R}$, we will sometimes use the notation $\varphi(t^{+})$ for $\lim_{s\downarrow t}\varphi(s)\coloneq\lim_{s>t,s\to t}\varphi(s)$ and the notation $\varphi(t^{-})$ for $\lim_{s\uparrow t}\varphi(s)\coloneq\lim_{s<t,s\to t}\varphi(s)$.
@@ 2 The adjoint state method: smooth case
This section studies the adjoint state method in the setting where the objective function $F$ is continuously differentiable, in the presence of equality and inequality constraints and possibly a continuum of minimizers. We focus here on the smooth case for simplicity and will expand on nonsmooth objectives in Section 4.
@@ 2.1 Preliminaries
where $\theta\in\mathbb{R}^{q}$. The objective function $F:\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}$ is locally Lipschitz, and the constraint mappings
@@
are continuously differentiable, the inequality being understood coordinate-wise. For any $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$, let $I(x,\theta)\subset\{1,\ldots,m\}$ be the set of active (inequality) constraints, i.e., for all $i\in I(x,\theta)$, $g_{i}(x,\theta)=0$. The Lagrangian of Problem (2.1), $\mathcal{L}\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\times\mathbb{R}^{q}\to\mathbb{R}$, writes
@@
Let $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $G(x,\theta)\leq 0,H(x,\theta)=0$. We say that the MFCQ holds at $(x,\theta)$, if the set
$\left\{\nabla_{\!x}h_{j}(x,\theta),j=1,\ldots,p\right\}$
is linearly independent and there exists a direction $d\in\mathbb{R}^{n}$ such that
$\nabla_{\!x}g_{i}(x,\theta)^{\!\mathsf{T}}d<0\text{ for all }i\in I(x,\theta)\text{ and }\operatorname{Jac}_{x}\!H(x,\theta)d=0.$
The MFCQ has an equivalent dual formulation. Indeed, the MFCQ is satisfied at $(x,\theta)$ if and only if
@@
If the MFCQ is satisfied at $(x,\theta)$, then the following holds for the joint Jacobians in $x,\theta$,
@@
$\implies\mu=0,\lambda_{i}=0,\forall\,i\in I(x,\theta).$ (2.3)
@@
    The objective function $F$ is continuously differentiable and is semialgebraic, i.e., its graph is described by polynomial inequalities and equalities. The constraint functions $G$ and $H$ are also semialgebraic.
@@ 2.2 The adjoint method
We propose the use of the adjoint formula from (1.2), without imposing any uniqueness requirement neither on the solutions to (2.1), nor on their associated dual/adjoint variables (Lagrange multipliers).
@@
In (2.7) and (2.6), one can recognize the operations of the adjoint state method proposed by [11, Equations (2.7) and (2.9)] with $m=0$ (no inequality constraint). The theoretical validity of the formula in [11], is based on the uniqueness and differentiability of both the primal and dual solutions, essentially requiring strong second order sufficiency conditions. The use of conservative fields enables us to consider the exact same formula as [11], without these strong conditions, allowing for possible nonsmooth or non-unique primal-dual solutions.
@@
where $D_{f}$ is a conservative field for $f$.
@@
    For general locally Lipschitz semialgebraic functions, like $f$, selections of conservative fields can be used for optimization. They carry sufficient variational information to serve as oracle for first-order methods with qualitative guaranties; see [8, 42]. This fact enables, for instance, parameter optimization for (2.1) through first-order methods by employing the adjoint state formula to efficiently compute a gradient-like oracle for $f$ at each iteration.
@@
    Conservative fields are compatible with differentiable programming paradigms such as algorithmic differentiation [8, 7]. This means that, in practice, ASM() can be combined with further operations in a larger numerical program, in a way which is compatible with algorithmic differentiation as implemented for example in Python libraries such as Pytorch.
@@
    The combination of [13, Proposition 2.1.2] and [8, Corollary 1] allows to obtain the following bound on Dini directional derivatives for any $\theta,d\in\mathbb{R}^{q}$:
@@
    This is reminiscent of the sensitivity analysis performed, for example, in [23, 33], where more precise results on directional derivatives are proposed.
@@
Set $n=2$ and $m=1$ with
$F(x,\theta)$ $=x_{1}x_{2},$ $G(x,\theta)=(x_{1},x_{2}-\theta,x_{2}+\theta)\!.^{\!\mathsf{T}}$
@@
2.3 Comparison to alternative methods
Throughout this section $\bar{\theta}\in\mathbb{R}^{q}$ denotes a fixed set of parameters. We compare ASM() with the following two widely used approaches that are based on differentiating the solution to get the sensitivities of $f$. The first approach, which we will simply call automatic differentiation, uses reverse-mode algorithmic differentiation to obtain the derivative of the solution $x(\bar{\theta})$ by differentiating through $\mathrm{solver_{p}}$. The second is implicit differentiation, where the implicit function theorem is applied to the critical condition $\nabla\mathcal{L}_{1}(\mathrm{solver_{pd}}(\bar{\theta}))=0$, with $\mathcal{L}_{1}\colon(x,\lambda,\mu)\mapsto\mathcal{L}(x,\lambda,\mu,\bar{\theta})$ where $(x,\lambda,\mu)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}$ and $\bar{\theta}$ is fixed.

### 22 Apr 2026 v2

@@ Abstract
We establish that nonconvex definable parametric optimization problems with possibly nonsmooth objectives, inequality constraints, conic constraint systems, and non-unique primal and dual solutions admit an adjoint state formula under a mere qualification condition. The adjoint construction yields a selection of a conservative field for the value function, providing a computable first-order object without requiring differentiation of the solution mapping. Through examples, we show that even in smooth problems, the formal adjoint construction fails without conservativity or definability, illustrating the relevance of these concepts to grasp theoretical aspects of the method. This work provides a tool which can be directly combined with existing primal-dual solvers for a wide range of parametric optimization problems.
Key words and phrases. Adjoint state method, conservative fields, nonsmooth optimization, conic programming, o-minimal geometry, parametric optimization, variational analysis.
2020 Mathematics Subject Classification. Primary 90C31; Secondary 03C64, 49J53, 65K05, 90C26, 90C30.
@@ 1 Introduction
Parametric optimization problems are central in optimization [10], control [37], and engineering [12]. They also arise naturally in modern machine learning, where one seeks to tune hyperparameters, architectures, regularization levels, or embedded optimization layers through the optimization of an induced value function; see [41, 4]. In these situations, the quantity of interest can often be recast in the parametric form
$\inf_{x\in\mathbb{R}^{n}}\left\{F(x,\theta):C(x,\theta)\in K\right\},$ (1.1)
where the parameter $\theta$ ranges in $\mathbb{R}^{q}$, $F\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}$ is locally Lipschitz, the constraint function $C\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{m}$ is continuously differentiable and $K$ is a closed convex cone in $\mathbb{R}^{m}$. For simplicity of exposition, we consider momentarily the following simpler instance
@@
where the constraint functions $G\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{m}$ and $H\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{p}$ are continuously differentiable, and the inequality is understood coordinate-wise. A central difficulty in optimizing the value function $f$ is to obtain useful first-order information at a reasonable computational cost. In principle, if one denotes by
@@
a selection of minimizers, the value writes $f(\theta)=F(x^{*}(\theta),\theta)$. Thus, if the mapping $\theta\mapsto x^{*}(\theta)$ were differentiable, the chain rule would give
@@
However, this approach raises both computational and conceptual difficulties. Computationally, differentiating through a solver is generally expensive, while conceptually it relies on strong regularity assumptions, such as those required by the implicit function theorem.
The adjoint state method, introduced by Céa [13], avoids computing the derivative of the solution mapping and does not resort to convex duality [44, 25]. Denoting $(x^{*}(\theta),\lambda^{*}(\theta),\mu^{*}(\theta))$ a solution of the KKT system of (1.2), the adjoint formula writes (see [13])
@@
Remarkably, at a negligible additional cost—see Remark 2.8()—associated with computing Lagrange multipliers, this formula does not involve the derivative of the primal-dual solution mapping. Yet, the classical validity analysis for the adjoint formula requires solutions of the KKT system to be unique and differentiable. To our knowledge, this requirement has not been relaxed; see the discussion in the related work section and Example 2.7().
@@
The objective of this work is to establish simple and versatile guarantees ensuring that the right-hand side of the adjoint formula (1.3) is a gradient-type quantity for the value function $f$.
For nonlinear programming problems (1.2), our results have an appealing form: assuming that the Mangasarian–Fromovitz constraint qualification holds and that all the mappings $F,G,H$ belong to a common o-minimal structure (for example, are semialgebraic), the formula yields a conservative field, and in particular a gradient almost everywhere. We actually show that a similar result holds in the much more general setting (1.1), under Robinson’s qualification and o-minimality.
@@
Our approach relies on two key ingredients: definability and conservative calculus. Definability provides a natural framework that encompasses most optimization and machine learning models [9, 18], while offering powerful stability and regularity properties that are essential for our analysis. Conservative calculus, on the other hand, supplies generalized differential objects suited to nonsmooth calculus and has already led to strong applications in optimization and learning; see [9, 18, 40]. The use of conservative fields for the value function is essential here: the formal application of the adjoint formula in (1.3) produces spurious outputs, even for smooth data $F,G,H$ resulting in a smooth value function $f$; see the very simple Example 2.7(). This is very common in nonsmooth differential calculus. Nonetheless, the outputs of the adjoint formula (1.3) form a conservative field; in particular, we obtain a gradient almost everywhere.
To conclude, we emphasize that, in the absence of definability, the adjoint formula may fail to carry variational meaning for the value function, even for locally Lipschitz, path differentiable and (point-wise) differentiable objective function with smooth constraints; see Section 5().
The adjoint formula and applications.
@@
Although primarily theoretical, this work provides a rigorous foundation for the practical use of the adjoint state method in settings where smoothness or uniqueness of solutions cannot be guaranteed. Such situations arise in a wide range of applications, including robust optimization [5], uncertainty quantification [1], optimal design of structures and systems [1], parametric decomposition [28, 23], PDE-constrained optimization [31], and discretized neural ODEs [14]. A detailed investigation of these applications is beyond the scope of the present paper. We also point out that the natural connections between conservativity, differentiable programming, and algorithmic differentiation make our approach amenable to modern computational frameworks [8, 9]. In Section 2.3(), we describe these connections with algorithmic differentiation as well as comparisons with alternatives to the adjoint formula in this context.
We discuss the computational cost of our adjoint approach from the perspective of differentiable programming, comparing it with other standard methods that are generally more costly; see Section 2.3(). To illustrate further this connection with applications, we provide a basic gradient-like algorithm based on the adjoint method in Section 4.4(). Our results show that the gradient method based on (1.3) as a first order oracle, exhibits a minimizing behavior: generically, it is attracted to critical points of the value function $f$ for vanishing step sizes; see Proposition 4.11(). This result can be extended to more advanced first order optimization algorithms; see Section 4.4().
Related work on the adjoint method.
The term “adjoint” comes from the adjoint-state in the literature on control and PDE-constrained optimization, in particular shape optimization. The method is closely related to the field of sensitivity analysis in constrained programming and to the notion of parametric optimality in nonsmooth analysis. Let us discuss references and connections to our work.
A general presentation of sensitivity analysis for shape optimization can be found in [50]. The adjoint formula in (1.3) was presented, to our knowledge, in the context of shape optimization [13]. In this context, the adjoint formula naturally occurs in infinite dimension and its main objects of interest are directional derivatives. Subsequent developments [19] were based on the connection with min-max optimization [16]. Recent extensions can be found in [20, 21, 22]. These developments share a lot of similarities with results developed for sensitivity analysis of constrained programs.
Sensitivity analysis is a broad field of research, which goes beyond shape analysis, a general account is given in [11] in infinite dimension. We focus the discussion on connections with the adjoint formula in (1.3) applied to problem (1.2), which initially occurred in a nonlinear programming context. Early results date back to Gauvin [27] with many subsequent extensions [26, 3, 33] see the review in [39] . A naive application of these results to our problem (1.3) would lead, at best, to only directional derivatives, rather than a directly computable gradient-like object for the value function, as provided by our approach. Moreover, it would require, when possible, solving costly optimization problems over the sets of primal and dual solutions. Directly obtaining a gradient using this sensitivity analysis approach requires strong uniqueness conditions as in the original argument of Céa. This is visible for example in [31] where uniqueness is assumed, in the context of PDE constrained optimization, and [23] where second order sufficiency conditions are assumed in the context of decomposition methods.
Beyond directional derivatives, let us mention the parametric optimality approach. [32] and [42, Theorem 10.13] established that the Clarke subdifferential of the value function is included in a convex set constructed with the subdifferential of the Lagrangian; see the discussion in Section 4.2(). However, as noted in the simple Example 2.7(), this inclusion is strict in general, showing that subdifferential theory is not adapted to this subject. A limitation that this paper overcomes through conservative gradients.
@@
Let $\ell,L\in\mathbb{N}$ be natural numbers. We denote by $\langle{\cdot,\cdot}\rangle$ the canonical Euclidean scalar product on $\mathbb{R}^{\ell}$ and $\lVert{\cdot}\rVert$ its associated norm. For $r>0$ and $x\in\mathbb{R}^{\ell}$, we denote by $\mathcal{B}_{c}(x,r)$ the closed ball centered at $x$ with radius $r$, and by $\mathcal{B}(x,r)$ its open counterpart. Let $A\subset\mathbb{R}^{\ell}$. We write $\operatorname*{cl}A$, $\operatorname*{int}A$, and $\operatorname*{conv\!}A$ for the closure, interior, and convex hull of $A$, respectively. The Jacobian and partial Jacobian of a vector-valued function $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}^{L}$ at $z=(x,y)\in\mathbb{R}^{\ell}$ are denoted by $\operatorname{Jac}\varphi(z)$ and $\operatorname{Jac}_{x}\!\varphi(x,y)$, respectively. In the same way, the gradient and partial gradient of a real-valued function $\varphi\colon\mathbb{R}^{\ell}\to\mathbb{R}$ at $z=(x,y)\in\mathbb{R}^{\ell}$ are written as $\nabla\varphi(z)$ and $\nabla_{\!x}\varphi(x,y)$, respectively. For $s,t\in\mathbb{R}$, we will sometimes use the notation $\varphi(t^{+})$ for $\lim_{s\downarrow t}\varphi(s)\coloneq\lim_{s>t,s\to t}\varphi(s)$ and the notation $\varphi(t^{-})$ for $\lim_{s\uparrow t}\varphi(s)\coloneq\lim_{s<t,s\to t}\varphi(s)$.
@@ 2 The adjoint state method: smooth case
This section studies the adjoint state method in the setting where the objective function $F$ is continuously differentiable, in the presence of equality and inequality constraints and possibly a continuum of minimizers. We focus here on the smooth case for simplicity and will expand on nonsmooth objectives in Section 4().
@@ 2.1 Preliminaries
where $\theta\in\mathbb{R}^{q}$. The objective function $F:\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}$ and the constraint mappings
@@
are all continuously differentiable, the inequality being understood coordinate-wise. For any $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$, let $I(x,\theta)\subset\{1,\ldots,m\}$ be the set of active (inequality) constraints, i.e., for all $i\in I(x,\theta)$, $g_{i}(x,\theta)=0$. The Lagrangian of Problem (2.1), $\mathcal{L}\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\times\mathbb{R}^{q}\to\mathbb{R}$, writes
@@
Let $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $G(x,\theta)\leq 0,H(x,\theta)=0$. The MFCQ at $(x,\theta)$ states that the following implication holds
@@
If the MFCQ is satisfied at $(x,\theta)$, then the following holds for the joint Jacobians in $(x,\theta)$, i.e,
@@
$\implies\mu=0,\lambda_{i}=0,\forall\,i\in I(x,\theta).$
@@
    The objective function $F$ is continuously differentiable and semialgebraic, i.e., its graph is described by polynomial inequalities and equalities. The constraint functions $G$ and $H$ are also continuously differentiable and semialgebraic.
@@ 2.2 The adjoint method
We propose the use of the adjoint formula from (1.3), without imposing any uniqueness or smoothness requirement neither on the solutions to (2.1), nor on their associated dual/adjoint variables (Lagrange multipliers).
@@
In (2.6) and (2.5), one can recognize the operations of the adjoint state method proposed by [13, Equations (2.7) and (2.9)] with $m=0$ (no inequality constraint). The theoretical validity of the formula in [13], is based on the uniqueness and differentiability of both the primal and dual solutions, essentially requiring strong second order sufficiency conditions. The use of conservative fields enables us to consider the exact same formula as [13], without these strong conditions, allowing for possible nonsmooth or non-unique primal-dual solutions.
@@
where $D_{f}$ is a conservative field for $f$. In particular, for any locally Lipschitz $\theta\colon\mathbb{R}\to\mathbb{R}^{q}$, we have
[…]
@@
The same result applies mutatis-mutandis to more general constraints of the form $C(x,\theta)\in K$ for a closed convex cone $K$. In this section, we focus on the simpler and more widespread form (1.2) for simplicity; see Section 4() for a general statement. This allows to cover conic programs such as paramteric second order cone programs (SOCP) and semidefinite programs (SDP).
We highlight the following important properties of conservative fields.
@@
    For general locally Lipschitz semialgebraic functions, like $f$, selections of conservative fields can be used for optimization. They carry sufficient variational information to serve as oracle for first-order methods with qualitative guaranties; see [9, 48]. This fact enables, for instance, parameter optimization for (2.1) through first-order methods by employing the adjoint state formula to efficiently compute a gradient-like oracle for $f$ at each iteration; see Section 4.4().
@@
    Conservative fields are compatible with differentiable programming paradigms such as algorithmic differentiation [9, 8]. This means that, in practice, ASM() can be combined with further operations in a larger numerical program, in a way which is compatible with algorithmic differentiation as implemented for example in Python libraries such as Pytorch; see Section 2.3().
@@
    The combination of [15, Proposition 2.1.2] and [9, Corollary 1] allows to obtain the following bound on Dini directional derivatives for any $\theta,d\in\mathbb{R}^{q}$:
@@
    This is reminiscent of the sensitivity analysis performed, for example, in [27, 11, 39], where more precise results on directional derivatives are proposed.
@@
Set $n=2$ and $m=3$ with
$F(x,\theta)$ $=x_{1}x_{2},$ $G(x,\theta)=(x_{1},x_{2}-\theta,x_{2}+\theta)\!\;^{\!\mathsf{T}}.$
@@
2.3 Algorithmic differentiation and alternative methods
Algorithmic differentiation implements the idea of differentiating numerical programs in an efficient and user friendly way [30]. This is one of the main building blocks of neural network training in machine learning. In its most simple form, a numerical program is simply a composition $P=f_{1}\circ f_{2}\circ\ldots\circ f_{L}$, where each $f_{i}$ is taken from a dictionary of basic numerical functions, compatible with the differentiable programming framework. The user specifies the structure of the composition with a numerical program, and algorithmic differentiation combines automatically derivative objects $d_{1},d_{2}\ldots,d_{L}$ associated to each $f_{i}$, to evaluate the derivative of $P$ numerically. Theorem 2.4() states that for any $\theta$, $\text{ASM}(\theta)\in D_{f}(\theta)$, where $f$ is the value function in (1.2), and $D_{f}$ is a conservative field for $f$. Conservative fields are compatible with algorithmic differentiation [9]. Therefore, an interpretation of Theorem 2.4() is that ASM can be used as a gradient surrogate $d$ for the value function $f$ in a differentiable programming framework. Let us discuss potential alternatives from the algorithmic differentiation literature [30].
For the remainder of this section $\bar{\theta}\in\mathbb{R}^{q}$ denotes a fixed set of parameters. We compare ASM() with the following two widely used approaches that are based on differentiating the solution to get the sensitivities of $f$. The first approach, which we will simply call automatic differentiation, uses reverse-mode algorithmic differentiation to obtain the derivative of the solution $x(\bar{\theta})$ by differentiating through $\mathrm{solver_{p}}$. The second is implicit differentiation, where the implicit function theorem is applied to the critical condition $\nabla\mathcal{L}_{1}(\mathrm{solver_{pd}}(\bar{\theta}))=0$, with $\mathcal{L}_{1}\colon(x,\lambda,\mu)\mapsto\mathcal{L}(x,\lambda,\mu,\bar{\theta})$ where $(x,\lambda,\mu)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}$ and $\bar{\theta}$ is fixed.